OpenStudy (anonymous):

A helicopter company charges $85 for the first kilometer of a trip and $5 for every kilometer after that. If the total cost of a trip was $365, how many kilometers were flown?

OpenStudy (anonymous):

57 including the first $85 dollar one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i got 57 kilometers were flown

OpenStudy (anonymous):

why is it 57 and not 56?

OpenStudy (anonymous):

your equation to solve is: 5x + 85 = 365

OpenStudy (anonymous):

because of the first kilometer

OpenStudy (anonymous):

@Dubba2000 is right

OpenStudy (anonymous):

yea... but where does that extra 1 come from

OpenStudy (anonymous):

from the 85 dollar kilometer

OpenStudy (anonymous):

85 - 1 kilometer (365 - 85)/5 = 56 56 + 1 = 57

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

why 85-1 ... im really confused about that part

OpenStudy (anonymous):

365-85=280 1 kilometer 280/5=56 56 kilometers 1+56+ 57 kilometers theres your answer

OpenStudy (anonymous):

ohhhhh okay got it thanks

OpenStudy (anonymous):

anytime